# Shape-specific fluctuations of an active colloidal interface

### v1

@@ V Discussion and summary
The scalings of $W_{\theta}$ in for $N^{\prime}<<N$ is presented in Fig. 8. We find that the scaling law of Eq. (14) is satisfied; the ballistic $\beta_{\theta}^{\prime}=1$ is retained whilst the roughness $\alpha_{\theta}^{\prime}\approx 1.22(03)$ is obtained. Note that this is opposed to the smoothness exponent obtained on the full CI. We also note that the specific choice of $N^{\prime}\in[5,12]$ used here, thus at most $\approx 4\%$ of the chain length, is substantially smaller than in Fig. 5. For any appreciably larger $N^{\prime}$ the scaling of (14) is not observed. These results thus further highlight the fact that the smoothness exponent found in Fig. 4 is a sole consequence of the CI; with distinct topology specific results vis-a-viz a flat surface.
Figure 8: Scaling of orientational fluctuations in locally flat region. A $W_{\theta}$ for different locally flat lengths $N^{\prime}$, with the dynamic $\beta_{\theta}^{\prime}$ labelled. The same scaling is plotted in the inset. B Roughness scaling for $\langle W_{\theta}^{\text{SS}}(N^{\prime})\rangle_{t}$. Here, $N=312$ and $\tau_{f}=0.2$, $\Lambda=0.2$, $\tilde{\Lambda}=0.01$.
.6 Effect of $\tau_{r}$
The effect of $\tau_{r}$ is not studied any further in the main text. For $\tau_{r}\gtrsim 20$ an identical FV scaling for $W_{h}$ is obtained, whereas $W_{\theta}$ only approximates the scaling law (15) - See Fig. 9. For $W_{h}$ one sees a crossover at $\tau_{r}$ for the angular fluctuations between a diffusive $\beta_{\theta}\approx 0.5$ and a deterministic $\beta_{\theta}\approx 1$. The effect of randomness prohibits a systems-spanning $l_{c}$. Nevertheless, a smoothness exponent with an identical scaling can be obtained.
For $\tau_{r}\lesssim 10$ there is no longer a propelling C-shape, and indeed other DSS are observed (see Fig. 7). The reason for this is that the C-shape DSS is sensitive to long-wavelength fluctuations when $D_{r}\gtrsim 1$, such that any fluctuations on the length scale $\frac{b^{2}D_{r}}{v_{s}}$ is greater than typical monomer lengths. A physically compelling reason on the other hand is that for equilibrated systems $\tau_{r}=\frac{8\pi\eta b^{3}}{k_{b}T}$, with $\tau_{t}=\frac{6\pi\eta b^{2}}{k_{b}T}$ ($\eta$ the viscosity of the medium), thus we will normally have $\tau_{r}<<\tau_{t}$, and hence $D_{r}$ can be safely ignored under this consideration. In addition the model in (2) ignores hydrodynamic advection.
Figure 9: Height and orientational fluctuations for finite $\tau_{r}$. A $W_{h}$ for different chain lengths NN, with the dynamic $\beta_{h}$ labelled. Inset: Scaling for $W_{h}^{\text{SS}}(N)$. B $W_{\theta}$ for different chain lengths NN, with the dynamic $\beta_{\theta}$ labelled. Inset: Scaling for $W_{\theta}^{\text{SS}}(N)$. C Crossover for $W_{\theta}$ in the presence of finite $\tau_{r}$ (vertical dashed black line). Here, $\tau_{f}=0.2$, $\Lambda=0.2$, $\tilde{\Lambda}=0.01$ and $\tau_{r}=70$.
.7 History-dependence
Figure 10: Fluctuations of model with chemical trails. Example of A $W_{h}$ and $W_{\theta}$ evolution, with the dynamic $\beta_{h}$ labelled. B Orientation profile of the C-shape, taken at different time points. C Dynamical evolution of $\theta_{i}$ for selected monomers ii.. Here, $D_{c}=1$, $t_{0}=1$, with $N=64$ and $T=10^{5}$.
We present the results for $W_{h}$ and $W_{\theta}$ in Fig. 10 (A). We find that the scalings of $W_{h}$ are reproduced. With the micellar trails, when one starts such a system with arbitrary initial conditions (e.g. random), the trails deposited via (31) give rise to a forward-backward symmetry breaking in the orientational response to the chemicals. Thus, the system breaks the initial symmetry imposed on it and further universally picks up the C-shape. In our system, we choose sufficiently random initial conditions such that this effect is seen; we then compute $\Delta h$ once the chain has a well-defined interface. For a system with $N=64$, averaged over $10$ realizations, we also find a crossover from the growth exponent $\beta_{h}^{(1)}=0.25$ to $\beta_{h}^{(2)}\approx 1.53$, similar to the model without chemical trails. We note that the steady-state value is not fully reached, as it is prohibitively long to simulate this region. For the same reason, we could not perform a rigorous systems-size scaling analysis; nevertheless, we conjecture that $\alpha_{h}$ would remain unchanged.
For the orientation fluctuations, though the qualitative form is reproduced, we observe that the exact values of $\beta_{\theta}$ are different from that found in Section IV.3. In particular, we see from Fig.10 (A) (yellow line) that $\beta_{\theta}$ is also super-ballistic, its growth comparable to $\beta_{h}$. This enhanced growth in the angular sector in this case can be appreciated via both the steady-state profile (Fig.10(B)) and the dynamical evolution (Fig.10(C)) of the orientations - these are to be compared with Fig.4 (C) and (D) respectively. In particular, we note that the time evolution contains higher order temporal derivatives, especially for monomers near the edges (e.g. $i=9,50,55$ in Fig.10(C) versus $i=311$ of Fig.4(D)), thus any fluctuations about the mean will deviate from a simple ballistic scaling. In addition, the steady-state angular profile is no longer monotonically varying across the chain (compare Fig.10(B) versus Fig.4(C)). Nevertheless the symmetry of $\theta_{i}$ across the chain leads us to conjecture that there will also be a smoothness exponent in this case.
Our conclusions of this section are, thus, that the addition of trail-mediated interactions only appreciably affects the growth of $W_{\theta}$, with the growth and roughness of $W_{h}$, along with smoothness of $W_{\theta}$ unchanged.
@@ Acknowledgments
We thank Professors ME Cates and M Muthukumar for useful discussions. AGS acknowledges funding from the DIA Fellowship from the Government of India. TB is supported through the Luxembourg National Research Fund (FNR), grant reference C22/MS/17186249.

### v2

@@ VI Summary and Discussions
The scalings of $W_{\theta}$ in for $N^{\prime}<<N$ is presented in Fig. 8. We find that the scaling law of Eq. (14) is satisfied; the ballistic $\beta_{\theta}^{\prime}=1$ is retained whilst the roughness $\alpha_{\theta}^{\prime}\approx 1.02\pm 02$ is obtained. Note that this is opposed to the smoothness exponent obtained on the full C-shape. We also note that the specific choice of $N^{\prime}\in[5,21]$ used here, thus at most $\approx 4\%$ of the chain length, is substantially smaller than in Fig. 5. For any appreciably larger $N^{\prime}$ the scaling of (14) is not observed. These results thus further highlight the fact that the smoothness exponent found in Fig. 3 is a sole consequence of the C-shape; with distinct topology specific results vis-a-viz a flat surface.
.7 Other DSS for large NN
It is to be noted also that the phase diagram in Fig. 1 being $N$ specific and only comparing ($\Lambda$, $\tilde{\Lambda}$) does exclude other DSS that arise from simulation.Examples of these are shown in Fig. 9. Four examples are shown, we label them as (i) “Alt-C”, (ii) “Random front”, (iii) “Stiff-C”, and (iv) “Pinched front” respectively. We note that DSS(ii) corresponds to the “Disordered” phase in Fig. 1. For each of these DSS, the chain propels in a deterministic direction along the x-axis; thus (8) can be computed in principle. However, none of these DSS satisfy all the criteria enumerated in Sec. III.1. For DSS(i)-(iv), criteria (i) 2, 3, and 4, (ii) 1 and 3, (iii) 1,3 and 4; and (iv) 2 and 3 are satisfied. Thus, the C-shape is the unique topology that satisfies all four criteria. The hypothetical fluctuations for $W_{\theta}$ for these can also in principle be computed, but we do not pursue it here.
@@ Acknowledgments
We thank Professors ME Cates and M Muthukumar for useful discussions. We also thank an anonymous referee for their feedback and suggestions, which led to an improvement in the presentation of our results. AGS acknowledges funding from the DIA Fellowship from the Government of India. TB is supported through the Luxembourg National Research Fund (FNR), grant reference C22/MS/17186249.